Given a transformation T:R^2→R^2 defined as T(x_1,x_2 )=(0,x_1-x_2). Find ker⁡(T) and R(T), range of T


1
Expert's answer
2021-09-02T07:25:17-0400
"T\begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}=\begin{bmatrix}
 0 & 0 \\
 1 & -1
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}"

"\begin{bmatrix}
 0 & 0 \\
 1 & -1
\end{bmatrix}\to\begin{bmatrix}
 1 & -1 \\
 0 & 0
\end{bmatrix}"

"x_1=x_2"

"\ker(T)=\{(x_1, x_2)\in\R^2|x_1=x_2\}"



Hence "\{(1,1)\}" is a basis of "\ker(T)."

The set "\{(1,0)\}" forms a basis for the range of "T, R(T)."


Rank theorem


"rank(T ) + nullity(T ) = dim(\R^2)"

"1+1=2"